# YAYI-UIE: A Chat-Enhanced Instruction Tuning Framework for Universal Information Extraction

### v2

@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated tremendous capabilities in solving a variety of natural language tasks and are equipped with strong generalization abilities. Therefore, Lu et al. Lu et al. (2022) first introduced the concept of universal IE to uniformly model various IE tasks. They also proposed a large-scale pre-trained universal IE model called UIE. However, UIE still requires model fine-tuning for different IE tasks and schemas, which leads to its poor performance on unseen data. Lou et al. Lou et al. (2023) proposed USM and designed three unified token-linking operations to decouple various IE tasks, but its training and inference processes suffer from inefficiency. Wang et al. Wang et al. (2023) developed an end-to-end unified information extraction framework InstructUIE based on instruction tuning, which utilizes descriptive instructions to enable LLMs to understand different IE tasks. Nevertheless, these existing methods are deficient in their IE capabilities for Chinese languages other than English.
In this paper, we propose YAYI-UIE, an end-to-end chat-enhanced instruction tuning framework for universal information extraction that supports both Chinese and English. Our framework consists of the following two instruction-tuning steps. The first step involves utilizing dialogue data to fine-tune a base large-scale language model for obtaining a chat model with common understanding abilities. In the second step, we focus on enhancing the chat model’s performance in IE tasks. To achieve this, we construct the largest and most comprehensive Chinese IE benchmark dataset and combined it with the existing English benchmark. The universal IE model is obtained by instruction-tuning the chat model using this combined dataset. Experimental results show that our proposed framework achieves state-of-the-art performance on Chinese datasets under supervised settings, and has remarkable generalization performance under zero-shot settings. Furthermore, the ablation study verifies the effectiveness of chat-based instruction tuning for the universal information extraction task.
The main contributions of our work are as follows:
@@ 2 Methodology
In this section, we describe the proposed chat-enhanced instruction tuning framework for universal information extraction (YAYI-UIE). We start with the task schema and the inference procedure of our universal information extraction framework. And then we introduce the design of the two-step instruction tuning, including instruction tuning for chat and instruction tuning for information extraction.
@@
  • Instruction is a natural language text sequence that includes three elements: task type, task option, and output format. It consists of a description of the task type (e.g., extract all entities from the following input) to specify the task; a description of the task option to restrict the range of the labels in the output; and a description of the desired format of the output.
@@ 2.1 Instruction Tuning for Chat
[…]
@@ 2.2 Instruction Tuning for IE
After training on the chat data, the chat model gains a fundamental understanding of open-world language and has been further enhanced in its Chinese language capabilities. In the second step of the proposed two-step instruction tuning, we adapt the model to the IE tasks via the IE instruction datasets and standardize the output format of the model. Moreover, we construct the most comprehensive Chinese IE instruction benchmark dataset to support the supervised fine-tuning for information extraction.
@@
Due to the lack of Chinese datasets in existing information extraction benchmarks, we collect 16 Chinese datasets for NER, RE, and EE tasks from diverse domains to build a comprehensive Chinese instruction benchmark, and then combine it with the existing English benchmark IE INSTRUCTIONS Wang et al. (2023).
Figure 3 gives an overview of the English and Chinese information extraction data for instruction tuning, which includes the distribution of the data across different tasks, domains, and languages. It covers more than 10 domains, such as general, finance, biology, and healthcare. Specifically, the built IE data covers various label types. For the NER task, the Chinese datasets cover 28 entity types (e.g., people, locations, organizations, body parts, medications), and the English datasets cover 130 entity types (e.g., animal, weapon, conference, book, etc). For the RE task, the Chinese datasets cover 232 relation types (e.g., purchasing capital, increasing holdings, restructuring, nationality, alias, relatives, subscribing to shares, transferring), and the English datasets cover 236 relation types (e.g., founded by, state or province of headquarters, an employee of, occupation, creator). For the EE task, the Chinese datasets cover 84 event types (e.g., winning a bid, executive changes, release), and the English datasets cover 45 event types (e.g., born, demonstrate, meet, end organization, divorce). Additionally, the Chinese and English portions in the built IE data account for 54% and 46% respectively.
@@
To enhance the generalization ability, we perform negative sampling on the labels of each instance during the training phrase. For input text $t$ containing $n$ types of labels $L=\{l_{1},l_{2},\cdots,l_{n}\}$, we randomly add several labels to $R$ that do not belong to $R$.
During the training process, we fine-tuned the chat LLM on the IE corpus to obtain the universal IE model:
[…]
@@ 3 Experiments
In this section, we conduct experiments under both supervised settings and zero-shot settings to evaluate the effectiveness of YAYI-UIE. For implementation, we choose Baichuan2-13B Yang et al. (2023) as the backbone model and perform the proposed chat-enhanced instruction tuning on it with the $10^{-5}$ learning rate. For the evaluation metrics, we adopt the F1 value to evaluate each dataset in NER, RE and EE tasks in a strict matching manner, and report the respective average F1 of the English datasets and the Chinese datasets on the three tasks. Specifically, we match the entity span and the entity type for NER, match the relation triple for RE, and match the event trigger and argument for EE respectively.
@@ 3.1 Experiments on Supervised Settings
We compare YAYI-UIE with three representative universal information extraction methods and comparable task-specific methods that have state-of-the-art results.
@@ 3.1.2 Baselines
  • BERT-base Kenton and Toutanova (2019) refers to task-specific supervised models with state-of-the-art results based on the pre-trained language model BERT, which has been extensively applied to various tasks as a text encoder.
@@ 3.1.3 Results
Table 1 gives the experimental results of the comparative methods and YAYI-UIE on 20 English NER datasets. As shown in the table, YAYI-UIE achieves a higher average F1 value than UIE and the task-specific baselines, and gains 0.86% F1 improvement compared to the BERT-base methods. The BERT-based methods require learning separate models for different tasks and schemas, thus it cannot leverage the correspondence between tasks. As InstructUIE models the common capabilities across different IE tasks, it achieves better performance than the BERT-based methods. Compared to the BERT-based methods, YAYI-UIE outperforms them on 11 out of 20 datasets. When compared to the strong baseline InstructUIE, YAYI-UIE performs slightly worse. The possible reason is the backbone model and training data of InstructUIE are both limited to English only, while the backbone and fine-tuning data of YAYI-UIE are primarily in Chinese, which may affect its capability on English datasets.
Table 2 gives the experimental results of the comparative methods and YAYI-UIE on 4 Chinese NER datasets. As the existing universal information extraction methods only support the English language, we compare YAYI-UIE to the strong BERT-based methods. In the table, YAYI-UIE achieves the highest average F1 value of 91.19%. For the CCKS 2017, CCKS 2018 and MSRA, YAYI-UIE is only off by less than 2% in F1 values, while for eCommerce, it achieves an improvement of 14.37%. The experimental results demonstrate that our model outperforms the comparative models on the Chinese NER task.
@@
Table 4 gives the experimental results of the comparative models and YAYI-UIE on 2 Chinese RE datasets. From the table, we can see that YAYI-UIE achieves the highest average F1 value, and gains 8.23% F1 improvement compared to the baselines. In general, the experimental results demonstrate the effectiveness of our YAYI-UIE for both English and Chinese RE tasks.
@@ 3.2 Experiments on Zero-shot Settings
To validate the zero-shot performance of YAYI-UIE, we conduct experiments on NER, RE and EE to compare YAYI-UIE with four strong baseline methods.
@@ 3.2.1 Datasets
3.2.2 Baselines
  • ZETT Kim et al. (2022) is a novel framework that extracts relation triplets from unstructured text. ZETT is particularly efficient in zero-shot tasks because the framework reduces triplet extraction to a template infilling task, and then generates triplets based on templates for both seen and unseen data.
  • ChatGLM Du et al. (2022) is an open-source, Chinese-English bilingual conversation language model based on the architecture of the general language model (GLM).
  • KnowLM Zhang et al. (2023) is an open-source and extensible knowledge graph extraction tool that can realize extraction tasks such as named entity recognition and relation extraction, and also support instruction compliance and self-define types and formats.
3.2.3 Results
Named Entity Recognition
Table 7 gives the zero-shot experimental results of the comparative models and YAYI-UIE on 5 unseen English NER datasets, and 3 unseen Chinese NER datasets. For the English NER task, compared to the GPT series model, YAYI-UIE falls some way short of Chatgpt’s results. Compared with the current SOTA model InstructUIE, our model outperforms all the datasets except Literature and Music, in which the Micro-F1 score ranges from 1.23% to 3.27%. Noted that USM achieves the sota on Literature, the reason may be that USM is trained on the same task corpus and tested on the label held out, while our model has never seen the task corpus. The overall performance in English still falls the Chatgpt, the reason may be that our English ability has decreased due to the addition of Chinese data training. However, for the Chinese NER task, our model performs much better than these two GPT series models. We can observe that YAYI-UIE outperforms the current sota model UIE in Micro-F1 score on all the datasets except weibo, ranging from 1.55% to 8.61%. And the overall performance in Chinese is the best.
@@ 3.2.3 Results
Table 9: Zero-shot performance on EE task, including 2 English datasets and 2 Chinese datasets.
Relation Extraction
Table 8 gives the zero-shot experimental results of the comparative models and YAYI-UIE on 2 unseen English RE datasets, and 3 unseen Chinese RE datasets. We can observe that our model achieves the sota on both English and Chinese datasets. For the English RE task, our model outperforms the current sota model InstructUIE in average Micro-F1 score, which performs over 5 points better on the Wiki-ZSL. Our model performs much better than GPT series models, and ChatGLM2-6b performs poorly because of the difficulty in structurally parsing its results. The performance on FewRel is not obviously due to the small size and insufficient learning of the model. For the Chinese RE task, our model performs much better than other models. Both Chatgpt and KnowLM perform poorly, especially with KnowLM completely unable to output correct results.
Event Extraction
Table 9 gives the zero-shot experimental results of the comparative models and YAYI-UIE on 1 unseen English EE dataset and 2 unseen Chinese EE datasets. The result shows that our model has better Chinese ability. When compared to the ChatGPT and UIE models, our model significantly outperforms in average Micro-F1 score on both the event trigger extraction and event argument extraction task. We can observe that YAYI-UIE outperforms the current SOTA model, especially on the dataset of FewFC, ranging from 18.66% to 31.05%. For the English EE task, including the Event Trigger and Event Argument, YAYI-UIE does not perform well compared to InstructUIE, while still outperforming the ChatGPT. It is worth mentioning that InstructUIE only has English capability, and our model has added a large amount of Chinese data in the training, which has reduced the English capability of the model to some extent. We think the huge difference in results may be due to this reason.
4 Ablation Study
@@ 5 Related Work
The advent of large language models (LLMs) has instigated a revolutionary paradigm shift within the field of natural language processing Guo et al. (2023); Qin et al. (2023); Bubeck et al. (2023). LLMs, such as LLaMA Touvron et al. (2023a, b), ChatGPT Ouyang et al. (2022) and GPT4 OpenAI (2023), have exhibited remarkable abilities across various applications. These LLMs undergo three primary training stages: pre-training, supervised fine-tuning (SFT), and reinforcement learning from human feedback (RLHF). During the pre-training phase, LLMs gain extensive skills and knowledge. However, they face a challenge in adhering to specific instructions. To mitigate this limitation, SFT is incorporated as a supplementary step. This process entails additional training of the LLM utilizing a dedicated annotated dataset that includes instructions and corresponding responses, augmenting its capabilities in accurately following instructions. While the process of SFT successfully makes LLMs follow user instructions, there remains potential for further enhancement in aligning with human preferences and values. RLHF, by incorporating human feedback into the training loop, serves as a pivotal mechanism for steering LLMs toward generating high-quality and harmless responses. In this paper, we propose a chat-enhanced instruction tuning framework for universal information extraction, which consists of two steps in the SFT process. We first utilize dialogue data to fine-tune a base LLM for obtaining instructions following abilities. We then construct the largest and most comprehensive Chinese IE benchmark dataset and combined it with the existing English benchmark to fine-tune the chat model to generate the target IE structure.
@@
Information extraction (IE) constitutes a longstanding field devoted to the automated extraction of diverse information structures from unstructured textual sources. Classic IE metrics Mengge et al. (2020); Wang et al. (2021); Qin et al. (2021); Wang et al. (2022a) necessitate the formulation of task-specific architectures and the training of dedicated models, which reveals limitations in the generalization ability of models across diverse IE tasks and imposes stringent demands for annotated data. Lu et al. Lu et al. (2022) have recently pioneered UIE by uniformly modeling IE tasks with a text-to-structure framework, which utilizes the encoder-decoder architecture to generate SEL containing specified types and spans. However, a notable limitation of UIE lies in its deficiency in transfer learning capabilities across diverse tasks or schemas. Lou et al. Lou et al. (2023) proposed USM by designing three directed token-linking operations to decouple task-specific IE tasks into two extraction abilities. This approach, however, engages in semantic matching for each word, resulting in a notable increase in both training and inference time. In the era of LLMs, there are several research applying LLMs to IE tasks. Wang et al. Wang et al. (2023) proposed InstructUIE by utilizing instructive guidance to direct LLMs toward the task, facilitating the generation of target structures. Unfortunately, this method is deficient in IE capabilities for Chinese languages other than English. In this paper, we propose an end-to-end framework YAYI-UIE for universal information extraction that supports both Chinese and English.
@@ 6 Conclusion
In this paper, we propose a chat-enhanced instruction tuning framework YAYI-UIE for universal information extraction, and build the most comprehensive Chinese IE instruction benchmark. The proposed framework consists of two instruction-tuning steps. It first utilizes dialogue data to fine-tune a base large-scale language model for obtaining common understanding abilities, and then utilizes the constructed Chinese IE benchmark dataset along with the existing English benchmark for IE instruction fine-tuning. Experimental results show that our proposed framework achieves state-of-the-art performance in Chinese while maintaining English language capabilities. Additionally, the ablation study verifies the effectiveness of chat-based instruction tuning for the universal information extraction task.

### v3

@@ Abstract
YAYI-UIE: A Chat-Enhanced Instruction Tuning Framework for Universal Information Extraction
[…]
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated tremendous capabilities in solving a variety of natural language tasks and are equipped with strong generalization abilities. Therefore, Lu et al. Lu et al. (2022) first introduced the concept of universal IE to uniformly model various IE tasks. They also proposed a large-scale pre-trained universal IE model called UIE. However, UIE still requires model fine-tuning for different downstream tasks, which leads to its poor performance on unseen data. Lou et al. Lou et al. (2023) proposed USM and designed three unified token-linking operations to decouple various IE tasks, but its training and inference processes suffer from inefficiency. Wang et al. Wang et al. (2023) developed an end-to-end unified information extraction framework InstructUIE based on instruction tuning, which utilizes descriptive instructions to enable LLMs to understand different IE tasks. Nevertheless, these existing methods are deficient in their IE capabilities for Chinese languages other than English.
In this paper, we propose YAYI-UIE, an end-to-end chat-enhanced instruction tuning framework for universal information extraction that supports both Chinese and English. Our framework consists of the following two instruction-tuning steps. The first step involves utilizing dialogue data to fine-tune a base LLM for obtaining a chat model with common understanding abilities. In the second step, we focus on enhancing the chat model’s performance in IE tasks. To achieve this, we construct the largest and most comprehensive Chinese IE benchmark dataset and combined it with the existing English benchmark. The universal IE model is obtained by instruction-tuning the chat model using this combined dataset.
@@ 2 Methodology
In this section, we describe the proposed chat-enhanced instruction tuning framework for universal information extraction (YAYI-UIE). We start with the task schema and the inference procedure of our universal information extraction framework. Then we introduce the design of the two-step instruction tuning, including instruction tuning for chat and information extraction respectively.
@@
  • Instruction is a natural language text sequence that includes three elements: task type, task option, and output format. It consists of a description of the task type to specify the task; a description of the task option to restrict the range of the labels in the output; and a description of the desired format of the output.
@@ 2.1 Instruction Tuning for Chat
$\textbf{LLM}_{chat}=\textit{SFT}\left(\textbf{LLM}_{base},~{}\mathcal{D}_{dialogue}\right)$ (2)
@@ 2.2 Instruction Tuning for IE
After training on the chat data, the chat model gains a fundamental understanding of open-world language and has been further enhanced in its Chinese language capabilities. In the second step of the proposed two-step instruction tuning, we adapt the model to the IE tasks via the IE instruction datasets and standardize the output format of the model. Moreover, we construct the most comprehensive Chinese IE instruction benchmark dataset to support the supervised fine-tuning for the IE tasks.
@@
Due to the lack of Chinese datasets in existing IE benchmarks, we collect 16 Chinese datasets for NER, RE, and EE tasks from diverse domains to build a comprehensive Chinese instruction benchmark, and then combine it with the existing English benchmark IE INSTRUCTIONS Wang et al. (2023).
Figure 3 gives an overview of the English and Chinese IE data for instruction tuning, which includes the distribution of the data across different tasks, domains, and languages. It covers more than 10 domains, such as general, finance, biology, and healthcare. Specifically, the built IE data covers various label types.
@@
To enhance the generalization ability, we perform negative sampling on the labels of each instance during the training phrase. For input text $t$ containing $n$ types of labels $L=\{l_{1},l_{2},\cdots,l_{n}\}$, we randomly add several labels to $L$ that do not belong to $L$. During the training process, we fine-tuned the chat LLM on the IE corpus to obtain the universal IE model:
$\textbf{LLM}_{ie}=\textit{SFT}\left(\textbf{LLM}_{chat},~{}\mathcal{D}_{ie}\right)$ (3)
@@ 3 Experiments
In this section, we conduct experiments under both supervised settings and zero-shot settings to evaluate the effectiveness of YAYI-UIE. For implementation, we choose Baichuan2-13B Yang et al. (2023) as the backbone model and perform the proposed chat-enhanced instruction tuning on it with the $10^{-5}$ learning rate. For the evaluation metrics, we adopt the F1 value to evaluate each dataset in NER, RE and EE tasks in a strict matching manner, and report the respective average F1 of the English datasets and the Chinese datasets on the three tasks.
@@ 3.1.2 Baselines
  • BERT-base Kenton and Toutanova (2019) refers to task-specific supervised models with state-of-the-art results based on the pre-trained language model BERT.
@@ 3.1.3 Results
Table 1 gives the experimental results of the comparative methods and YAYI-UIE on 20 English NER datasets. As shown in the table, YAYI-UIE achieves a higher average F1 value than UIE and BERT-base methods. When compared to the strong baseline InstructUIE, YAYI-UIE performs slightly worse. The possible reason is the backbone model and training data of InstructUIE are both limited to English only, while the backbone and fine-tuning data of YAYI-UIE are primarily in Chinese, which may affect its capability on English datasets.
Table 2 gives the experimental results of the comparative methods and YAYI-UIE on 4 Chinese NER datasets. As the existing universal information extraction methods only support the English language, we compare YAYI-UIE to the strong BERT-based methods. In the table, YAYI-UIE achieves the highest average F1 value of 91.19%. For the CCKS 2017, CCKS 2018 and MSRA, YAYI-UIE is only off by less than 2% in F1 values, while for eCommerce, it achieves an improvement of 14.37%. The experimental results show that our model outperforms the baselines on the Chinese NER task.
@@
Table 4 gives the experimental results of the comparative models and YAYI-UIE on 2 Chinese RE datasets. From the table, we can see that YAYI-UIE achieves the highest average F1 value, and gains 8.6% F1 improvement compared to the baselines. In general, the experimental results demonstrate the effectiveness of our YAYI-UIE for both English and Chinese RE tasks.
@@ 3.2.1 Datasets
Table 9: Zero-shot performance on EE task, including 1 English datasets and 2 Chinese datasets.
@@
3.2.2 Baselines
  • ZETT Kim et al. (2022) is a framework that extracts relation triplets from unstructured text.
  • ChatGLM Du et al. (2022) is an open-source, Chinese-English bilingual conversation language model.
  • KnowLM Zhang et al. (2023) is an open-source and extensible knowledge graph extraction tool that can extract entities and relations.
3.2.3 Results
Named Entity Recognition
Table 7 gives the zero-shot experimental results of the comparative models and YAYI-UIE on 5 unseen English NER datasets and 3 unseen Chinese NER datasets. For the English NER task, YAYI-UIE outperforms several strong baselines except for ChatGPT. For the Chinese NER task, YAYI-UIE achieves the highest average F1 score.
Relation Extraction
Table 8 gives the zero-shot experimental results of the comparative models and YAYI-UIE on 2 unseen English RE datasets, and 3 unseen Chinese RE datasets. We can observe that YAYI-UIE achieves the SOTA on both English and Chinese datasets. For the English RE task, YAYI-UIE outperforms the best comparative model InstructUIE in average F1 score by 1.2%. The performance on FewRel is not obviously due to the small size and insufficient learning of the model. For the Chinese RE task, our proposed model performs much better than the baselines.
Event Extraction
Table 9 gives the zero-shot experimental results of the comparative models and YAYI-UIE on 1 unseen English EE dataset and 2 unseen Chinese EE datasets. The result shows that YAYI-UIE achieves the SOTA performance for Chinese EE task, and also comparable performance for English EE task. It is worth mentioning that InstructUIE only has English capability, and our model has added a large amount of Chinese data in the training, which has reduced the English capability of the model to some extent.
4 Ablation Study
@@ 5 Related Work
The advent of large language models (LLMs) has instigated a revolutionary paradigm shift within the field of natural language processing Guo et al. (2023); Qin et al. (2023); Bubeck et al. (2023). LLMs, such as LLaMA Touvron et al. (2023a, b), ChatGPT Ouyang et al. (2022) and GPT4 OpenAI (2023), have exhibited remarkable abilities across various applications. These LLMs undergo three primary training stages: pre-training, supervised fine-tuning (SFT), and reinforcement learning from human feedback (RLHF). During the pre-training phase, LLMs gain extensive skills and knowledge. However, they face a challenge in adhering to specific instructions. To mitigate this limitation, SFT is incorporated as a supplementary step. This process entails additional training of the LLM utilizing a dedicated annotated dataset that includes instructions and corresponding responses, augmenting its capabilities in accurately following instructions. RLHF, by incorporating human feedback into the training loop, serves as a pivotal mechanism for steering LLMs toward generating high-quality and harmless responses. In this paper, we propose a chat-enhanced instruction tuning framework for universal information extraction, which consists of two steps in the SFT process. We first utilize dialogue data to fine-tune a base LLM for obtaining instructions following abilities, and then fine-tune the chat model to generate the target IE structure.
@@
Information extraction constitutes a longstanding field devoted to the automated extraction of diverse information structures from unstructured textual sources. Classic IE methods Mengge et al. (2020); Wang et al. (2021); Qin et al. (2021); Wang et al. (2022a) necessitate the formulation of task-specific architectures and the training of dedicated models, which reveals limitations in the generalization ability of models across diverse IE tasks and imposes stringent demands for annotated data. Lu et al. Lu et al. (2022) have recently pioneered UIE by uniformly modeling IE tasks with a text-to-structure framework. However, a notable limitation of UIE lies in its deficiency in transferring learning capabilities across diverse tasks or schemas. Lou et al. Lou et al. (2023) proposed USM by designing three directed token-linking operations to decouple task-specific IE tasks into two extraction abilities, resulting in a notable increase in both training and inference time. Wang et al. Wang et al. (2023) proposed InstructUIE by utilizing instructive guidance to direct LLMs toward the task, facilitating the generation of target structures. Unfortunately, this method is deficient in IE capabilities for Chinese languages other than English. In this paper, we propose an end-to-end framework YAYI-UIE for universal information extraction that supports both Chinese and English.
@@ 6 Conclusion
In this paper, we propose a chat-enhanced instruction tuning framework YAYI-UIE for universal information extraction, and build the most comprehensive Chinese IE instruction benchmark. The proposed framework consists of two instruction-tuning steps. It first utilizes dialogue data to fine-tune a base LLM for obtaining common understanding abilities, and then utilizes the constructed Chinese IE benchmark dataset along with the existing English benchmark for IE instruction fine-tuning. Experimental results show that our proposed framework achieves state-of-the-art performance in Chinese while maintaining English language capabilities.